OpenStudy (anonymous):

Question 3 Make sure you read the directions for each part carefully. This system of equations will be used in all three parts of this question. 2x+3y=6 x-2y=10 Part 1: Solve the system by graphing. Draw the graphs on paper and indicate their intersection point. Type the ordered pair solution here as your answer to Part 1. Part 2: Solve the system using the substitution method. Show all work here and indicate the solution for the system as an ordered pair. Part 3: Solve the system using the addition method. Show all work here and indicate the solution for the system as an ordered pair

OpenStudy (anonymous):

Part1: just identify 2 points for each equation (line) & look for point of intersection; Part3: re-write it as: 2x+3y=6 -2x+4y=-20 (multiplied 2nd eq by (-2)) add together - x will cancel out, solve it for y, then - >x Part2: from the equation: x-2y=10 x=2y+10 put the value of x in second equation & solve it for y; then - >x

OpenStudy (anonymous):

answers should be all the same: x=6, y=-2 Problems?

OpenStudy (anonymous):

no